Show that if W consist of these vectors (a, b, c)€R³ for which a=2b then W is subspace of R³

1
Expert's answer
2022-05-12T04:17:55-0400

"W" consists of vectors "(2b,b,c)", where "b,c\in{\mathbb{R}}". We will use a subspace criterion to check that "W" is a subspace of "{\mathbb{R}}^3". Namely, we have to check the following:

  1. "0\in W".
  2. "u+v\in W" for all vectors "u,v\in W".
  3. "\alpha u\in W" for any "\alpha\in{\mathbb{R}}" and for an arbitrary vector "u\in W".

It is obvious that "0\in W". Take two arbitrary vectors "u=(2b,b,c)" and "v=(2d,d,f)" from "W" with "b,c,d,f\in{\mathbb{R}}". Then, "u+v=(2(b+d),b+d,c+f)". As we can see, "u+v\in W". Consider arbitrary vector "u=(2b,b,c)" from "W" with "b,c\in{\mathbb{R}}" . Then, "\alpha u=(2\alpha b,\alpha b,c)\in W" for all "\alpha\in{\mathbb{R}}". Thus, we checked that "W" is a subspace of "\mathbb{R}^3".